# A Mixed-Methods Study of Open-Source Software Maintainers On Vulnerability Management and Platform Security Features

### v1

@@ 1 Introduction
Given the pervasiveness of OSS, OSS software vulnerabilities deep within the software supply chain can have a crippling effect on society [32, 28, 36]. At the forefront of defending against these vulnerabilities are OSS maintainers who oversee those projects, where many may not have experience handling a security incident [90]. OSS platforms such as GitHub offer various platform security features (e.g., dependency management [44]), but these features have been underutilized [8, 10, 9]. With the many responsibilities of being a maintainer, some have felt burned out [57] or left their project completely [54]. The recent xz-utils incident [70] in which an XZ Utils maintainer, who joined the project two years earlier, maliciously introduced a vulnerability in the project further highlights the challenges existing maintainers face in vulnerability management and mistrust in the software supply chain. While several studies focus on the perspectives of vulnerability reporters [2, 3, 51, 65, 34, 96, 61] and the security practices of OSS contributors [49, 90], there exist knowledge gaps concerning the perspectives of OSS maintainers, especially those who own previously vulnerable projects.
@@
We conduct two studies, described in Section 3, to address our RQs: a survey ($n_{1}=80$) to list current practices, general vulnerability management challenges, platform security feature challenges, platform security feature barriers, and wanted features or improvements, and conduct semi-structured interviews ($n_{2}=22$) study to contextualize overall survey results.
@@ 3.1.2 Survey details
For each question, we informed that study participants should list all tooling and factors they use when conducting vulnerability management. We use an open-ended listing approach for free-response questions called free-listing [13], common in research when a domain is understudied, to elicit a full breadth of responses. Next, we asked participants to self-report OSS maintenance and industry experience, if they have a security background, project funding, and how often their vulnerability management process is reviewed. We ended with demographic questions to understand our population. The full outline and questions asked can be found in Appendix A.
@@ 3.1.3 Data analysis
We analyzed listing survey responses with exploratory open-coding [80]. Two researchers independently coded batches of ten responses at a time; further, resolving differences and updating the codebook after each batch. Because the survey lent itself to multiple listings per category, we were able to create an initial codebook from 27/80 (33.8%) respondents, resulting in 23/37 (62.2%) factors in Table 2. In particular, two researchers analyzed the same 34 participant responses by engaging in open coding [22] and discussing the initial emerging themes, meeting five times. Both researchers then independently coded the remaining 46 responses in batches of 6 and met frequently to discuss findings and reach a consensus, resulting in 14 emerging codes after the initial analysis. The final 11 responses were received after the final codebook was established, containing 37 codes across categories in Table 2.
@@ 3.2.1 Participant recruitment and piloting
Interviewees were a subset of the initial Listing survey study described in Section 3.1. If respondents were interested in participating in the interview study, they were directed to a Calendly [1] space where they could join a publicly available Zoom link during specified time ranges and, if so, provide the GitHub project they oversee. This also allowed us to access their corresponding non-identifiable demographic information and background as they were prompted to select if they were interested in participating in the interview study. As a result, 35 agreed to be interviewed. After reaching out using publicly listed contact information from GitHub projects, 25 responded, 22 of whom confirmed that they were available for an interview during the time ranges we advertised.
We conducted pilots with three OSS project maintainers to test our semi-structured interview protocol, which can be found in Appendix B. We revised our interview questions based on the interviewees’ quality of responses to narrow our research focus. Details about participants regarding project information and interviews can be found in Appendix C.
@@ 3.2.2 Interview details and structure
For reporting, we group the interview into six sections, each consisting of 1-2 opening questions, corresponding follow-up questions, and circling back to earlier ideas when applicable. Before the main interview portion, we introduced our institution affiliations, an overall project overview, and our motivations. We went over how we only intended to keep audio, gathered the participant’s consent for recording, and enabled closed captions and transcription. We outline our semi-structured interview structure below and in Figure 1.
Figure 1: Illustration of the flow of topics of the semi-structured interview protocol. In each section, participants were presented with general questions and corresponding follow-ups but were free to diverge or skip questions entirely.
0. Project maintainer duties The interview opens with a general question section about the project and our participant’s relation to it. This section is intended both to ease nervous participants into the interview as well as establish some initial context to later combine with actual repository data.
1. Current vulnerability management practices This section investigates vulnerability management processes in the project. In particular, we were interested in structures that are not easily visible, such as who is involved in handling vulnerabilities, how vulnerabilities are actually handled, and participant perceptions of why current approaches are appropriate for them. We report these results in Section 5.1.
2. Challenges with vulnerability management This section investigates vulnerability management challenges in the project. In particular, we were interested in exploring past vulnerability management challenges encountered by our participants, e.g., through experiences, and how they have dealt with such challenges. We report these results in Section 5.2.
3. Challenges with platform security features This section investigates platform security feature-specific challenges, i.e., challenges with tooling. We were interested in what overhead participants face with the platform security features they use and how such challenges affect their overall vulnerability management practices. We report these results in Section 5.3.
4. Barriers to adopting platform security features This section investigates platform security feature-specific barriers, i.e., what prevents the adoption of such tooling. We were interested in why participants refrain from using built-in security features and their perceptions of what such features have to, or not, offer. We report these results in Section 5.4.
5. Opportunities for improvement and support This section aims to elicit participants’ wants and needs for improving current platform security features and vulnerability management practices in general. We also ask for advice, if any, that would benefit other OSS maintainers looking to make their project “secure.” We report these results in Section 5.5.
@@ 5 Results
In the following section, we report and discuss the results for 22 semi-structured interviews with open-source maintainers. In our reporting, we adhere to the structure of the interview guide shown in Figure 1 from Section 3.2. We report quotes as transcribed with minor grammatical corrections and omissions. e.g., using “[…]”, for readability.
@@
Factors and tooling Description RW L I
Current practices Email or mailing list A private email or mailing list for vulnerabilities. [90, 58] 50 11
Proactive disclosure Established disclosure process after a vulnerability patch. [90] 46 7
Security policy Instructs users how to report vulnerabilities. [90] 37 10
GitHub private reporting A built-in feature for privately reporting vulnerabilities. 20 9
Automation tooling E.g., Dependabot, code scanning, secret scanning. [76, 30, 48, 71, 4, 69] 16 12
GitHub Issues Vulnerabilities are reported publicly with GitHub Issues. [24] 9 6
External tooling Utilizes tooling or reports outside of GitHub. [48, 90, 35] 3 2
Ignores vulnerabilties Ignores vulnerability reports or does nothing in response. 2
General challenges Supply chain trust E.g., waiting for upstream dependencies to deploy a fix. [5, 90, 26, 75, 39] 36 6
Lack of understanding E.g., knowledge gaps, how to start developing a patch. [5, 7] 35 12
Lack of time Balancing vulnerability priority with other OSS tasks. [57, 90, 7] 25 9
Lack of resources E.g., “I do almost everything myself” (P4). [90, 20] 23 11
CVE relationships Negative experiences with CVE-assigned vulnerabilities. 8 9
Lack of procedure Maintainers are not sure how to deal with vulnerabilities. [90] 7 5
Disclosure coordination How to reach out to all users and what to disclose. [5] 7 4
Negative attitudes Past negative interactions with reporters or companies. [67, 78, 31, 54, 57, 29] 6 9
PSF challenges Not enough automation More automation needed in PSF functionality. [56, 18] 39 11
Too much noise E.g., too many false positives. [90, 89, 48, 69, 92, 91, 17] 24 10
Vulnerability scoring E.g., having a hard time calculating CVSS scores. 21 9
CI processes are missing GitHub Actions & tests are not available in private forks. 8 7
Broken tests & builds Failing builds & tests after vulnerability patch s applied. [94, 79, 69, 17] 6 7
Cluttered notifications E.g., vulnerable dependencies presentation is cluttered. 4 4
Too much manual setup E.g., manually enabling PSFs on all projects. 3 3
PSF barriers Lack of awareness E.g., not knowing about private vulnerability reporting. [9] 33 12
Complex to set up or use PSFs are too complex for maintainers with limited time. 26 7
They are unnecessary Perception that using PSFs is not needed. 23 6
Reputation concerns Past vulnerabilities reflects a negative project reputation. 11 4
Bad UI presentation Hard to find or hidden, “second-class features” (P10). [27, 55, 35] 8 6
Lack of motivation Feeling burned out from other OSS tasks. [78, 38, 64, 6, 57] 5 4
Not sure what they do Unsure of PSFs’ functionality, benefits, etc. 4 2
Maintainer wants Assisted triaging Automation for impact analysis, patch development, etc. [56, 18, 91] 39 9
Assisted PSF setup More guidance for the PSF setup process. 37 10
Security-specific funding Opportunities for funding OSS security efforts. [90] 32 7
User-friendly resources Tailored for those without a security background. 17 8
Gamification for projects E.g., a green shield for projects with proper PSFs set up. [72, 86, 26] 13 6
PSF checklist To-do list for enabling easily configurable PSFs. 15 3
Nudges or reminders Nudges reminding maintainers to consider using PSFs. 12 5
@@
Table 2: 37 factors and tooling codes, shortened to preserve space, identified from the listing survey study and interview study divided into five categories as a result of software vulnerability management efforts. Acronyms listed represent, respectively: platform security features (PSF), existing related work (RW), the number of listing (L) survey study mentions ($n_{1}=80$), and the number of interview (I) study mentions ($n_{2}=22$). The extended table without omitted codes can be found in Appendix D.
@@ 5.1 Current practices used for vulnerability management
Emails and external tooling or reports Using a form of a security contact email or mailing list consisting of project maintainers was the most listed form of current vulnerability management practices ($L=50$, $I=11$). This is consistent with a finding from Wermke et al., where authors identified that having a security contact point was the commonly mentioned aspect of a security policy [90].
@@
Proactive disclosure process Having an established disclosure process after patching vulnerabilities consisting of project maintainers was the second-highest form of current vulnerability management practices ($L=46$, $I=7$). This demonstrates a sense of responsibility within the OSS community since it ensures affected dependent client projects within reach are notified to upgrade [71].
@@
Security policies The next most listed current practice used is having a project security policy ($L=37,I=10$). OSS security policies have mostly been used to inform reporters how to properly communicate vulnerabilities to maintainers via security contact [90], but they are generally underused [8].
@@
Private vulnerability reporting Another GitHub platform security feature listing participants mentioned frequently is private vulnerability reporting ($L=20,I=9$). This particular feature allows contributors to privately report vulnerabilities within the GitHub platform. OSS maintainers can then review a report with functionalities to update its severity, invite others to develop a fix, and decide whether or not to request a CVE.
@@
Automation tooling Another method of current practices listing and interviewees mention is the usage of automated dependency analysis tooling and code scanning $(L=16,I=12)$. Dependency analysis provides alerts of library upgrades, e.g., patched versions, while code scanning can reveal vulnerabilities, e.g., secrets, as a result of running static analysis.
@@
“The first one that I use quite often is Renovate, that is a tool in Github easily available where you can configure: I want this and this upgraded like that, and you can have all kinds of settings and then it’s automatically gives you a pull request […] with a dependency update and automatically, the test pipeline fires.” (P1).
@@
GitHub Issues Some listing participants and interviewees indicate using GitHub issues for reporting vulnerabilities ($L=9,I=6$). When a GitHub Issue is submitted, it is publicly accessible in the Issues tab located at the top of the project landing page.
@@
Ignoring vulnerabilties Two listing participants mention how they ignore vulnerabilities. One attributed this behavior to a lack of motivation for setting up any security features or tooling and indicated that “financial support or sponsorship” would be most beneficial for them to improve their vulnerability management practices. The other respondent listed “unless the project is extremely sensitive - avoid implementing any policy” and expressed the perception that having security features enabled for projects is unnecessary. Though this respondent listed that they were interested in participating in our interview study, they did not respond with any contact information or reach out to us directly.
@@ 5.2 Challenges with vulnerability management
All interviewees described wearing many hats when it comes to OSS maintenance, some of who are lone maintainers ($I=11$). General vulnerability management challenges OSS maintainers face range from trusting the software supply chain to lack of time and resources to issues with CVEs. Participants also describe how a lack of standardized vulnerability handling procedures and implementing proper coordinated disclosure are prominent challenges. Some bring up the idea of having “imposter syndrome” when receiving vulnerability reports, while others share experiences of negative encounters with reporters and companies.
Supply chain trustworthiness Supply chain trust was the most listed challenge with vulnerability management ($L=36,I=6$). Use cases and duties related to the software supply chain in OSS include adopting upstream dependencies, tracking the status of dependencies, and updating dependencies in a timely manner, e.g., when a vulnerability is patched.
@@
Lack of understanding The second most listed challenge for conducting vulnerability management is a lack of understanding ($L=35,I=12$). Not being able to effectively understand reported vulnerabilities can lead to delays in patching and disclosure to affected dependent client projects.
@@
Lack of time and resources The next most listed software vulnerability management challenges are the lack of time ($L=25,I=9$) and resources ($L=23,I=11$). The time OSS maintainers spend on developing and managing projects is limited, e.g., hobbyists. Further, resources are scarce, e.g., large-scale software ecosystems like Python Package Index and NPM are primarily made up of projects with a single maintainer [15, 16].
@@
Negative relationships with CVEs Participants also expressed bad relationships with CVEs and the overall CVE process as a challenge when conducting vulnerability management ($L=8,I=9$). CVEs play an important role in the overall disclosure process and help promote transparency in the OSS ecosystem, e.g., upgrade notifications.
@@
Lack of procedures and coordinated disclosure Though established procedures can help alleviate the receive-to-resolve timeline, the lack of procedures is a challenge OSS maintainers also face ($L=7,I=5$). Further, participants also express challenges with being able to effectively reach affected dependent client projects and users ($L=7,I=4$).
@@
Negative attitudes The least listed challenge of vulnerability management is negative attitudes from reporters ($L=6,L=9$). The OSS ecosystem is human-centered, e.g., through community and collaboration, and is not a stranger to toxic interactions [78, 67]. One interviewee described experiences with companies demanding pentests. Participants mention the argumentative nature with vulnerability reporters ($L=3,I=7$) and those with sole intentions to make money ($L=3,I=5$).

### v2

@@ 1 Introduction
Given the pervasiveness of OSS, OSS software vulnerabilities deep within the software supply chain can have a crippling effect on society [37, 33, 42]. At the forefront of defending against these vulnerabilities are OSS maintainers who oversee those projects, where many may not have experience handling a security incident [101]. OSS platforms such as GitHub offer various platform security features (e.g., dependency management [52]), but these features have been underutilized [11, 14, 12, 15]. With the many responsibilities of being a maintainer, some have felt burned out [67] or left their project completely [63]. The recent xz-utils incident [80] in which an XZ Utils maintainer, who joined the project two years earlier, maliciously introduced a vulnerability in the project further highlights the challenges existing maintainers face in vulnerability management and mistrust in the software supply chain. While several studies focus on the perspectives of vulnerability reporters [4, 6, 60, 75, 39, 108, 71] and the security practices of OSS contributors [57, 101], there exist knowledge gaps concerning the perspectives of OSS maintainers, especially those who own previously vulnerable projects.
@@
We conduct two studies, described in Section 3, to address our RQs: a survey ($n_{1}=80$) to list current practices, general vulnerability management challenges, platform security feature challenges, platform security feature barriers, and wanted features or improvements, and conduct semi-structured interviews ($n_{2}=22$) study to contextualize overall survey results. Our primary contributions are as follows:
  • We are the first to investigate vulnerability management challenges that OSS maintainers, whose projects have a history of patched vulnerabilities, face regarding platform security features using GitHub Advisory Database.
  • We confirm previously discovered challenges, and further unveil novel challenges and barriers, OSS maintainers face when using current platform security features. Further, we systematically determine the breadth of security tooling and practices OSS maintainers use.
  • We provide an in-depth discussion, including suggestions and actionable items, which have corresponding insights to improve the OSS maintainer security processes, and have made our questionnaires for both studies publicly available [2] to facilitate future research.
@@ 3.1.2 Survey details
For each question, we informed that study participants should list all tooling and factors they use when conducting vulnerability management. We use an open-ended listing approach for free-response questions called free-listing [18], common in research when a domain is understudied, to elicit a full breadth of responses. Next, we asked participants to self-report OSS maintenance and industry experience, if they have a security background, project funding, and how often their vulnerability management process is reviewed. We ended with demographic questions to understand our population. The full outline and questions asked can be found in our artifact [2].
Prior work has not systematically determined the breadth of security tooling and practices OSS maintainers use. Instead, it has studied specific tooling in specific contexts, e.g., dependency tooling [8, 56, 79], static code analysis [65, 100, 105], and what OSS stakeholders implement for incident-handling [101, 11], while we focus on OSS maintainers since they are the primary point-of-contact for handling vulnerabilities. Prior work informs our survey design, which explores the breadth of approaches, challenges, and tooling GitHub OSS maintainers use to improve their projects’ security, which is currently understudied; our survey additionally draws from the Getting started GitHub security features guide [49] and established initiatives that inform OSS maintainers on how to approach vulnerability management, e.g., guides from OpenSSF [91].
@@ 3.1.3 Data analysis
We analyzed listing survey responses with exploratory open-coding [90]. Two researchers independently coded batches of ten responses at a time; further, resolving differences and updating the codebook after each batch. Because the survey lent itself to multiple listings per category, we were able to create an initial codebook from 27/80 (33.8%) respondents, resulting in 23/37 (62.2%) factors in Table 2. In particular, two researchers analyzed the same 34 participant responses by engaging in open coding [27] and discussing the initial emerging themes, meeting five times. Both researchers then independently coded the remaining 46 responses in batches of 6 and met frequently to discuss findings and reach a consensus, resulting in 14 emerging codes after the initial analysis. The final 11 responses were received after the final codebook was established, i.e., coders met synchronously to fit them within the appropriate existing codes,containing 37 codes in Table 2.
@@ 3.2.1 Participant recruitment and piloting
Interviewees were a subset of the initial Listing survey study described in Section 3.1. If respondents were interested in participating in the interview study, they were directed to a Calendly [1] space where they could join a publicly available Zoom link during specified time ranges and, if so, provide the GitHub project they oversee. Of the 35/80 who responded Yes, 19/35 scheduled a Calendly meeting. Using emails for recruitment, we contacted the other 16/35: 6 scheduled a meeting, and 10 did not respond. Of the 25 who scheduled a meeting, 3 were no-shows, resulting in 22 interviews.
We conducted pilots with three OSS project maintainers to test our semi-structured interview protocol, which can be found in our artifact [2]. We revised our interview questions based on the interviewees’ quality of responses to narrow our research focus. Details about participants regarding project information and interviews can also be found in our artifact [2].
@@ 3.2.2 Interview details and structure
For reporting, we group the interview into six sections, each consisting of 1-2 opening questions, corresponding follow-up questions, and circling back to earlier ideas when applicable. Before the main interview portion, we introduced our institution affiliations, an overall project overview, and our motivations. We went over how we only intended to keep audio, gathered consent for recording, and enabled closed captions and transcription. We proceeded as follows:
0. Project maintainer duties This section is intended to ease nervous participants into the interview and establish initial context to later combine with actual repository data.
1. Current vulnerability management practices This section explores structures that are not easily visible, such as who is involved in handling vulnerabilities, how vulnerabilities are actually handled, and participant perceptions of why approaches are appropriate for them (Section 5.1).
2. Challenges with vulnerability management This section explores past vulnerability management challenges encountered by our participants, e.g., through experiences, and how they have dealt with such challenges (Section 5.2).
3. Challenges with platform security features This section explores what overhead participants face with the PSFs they use and how such challenges affect their overall vulnerability management practices (Section 5.3).
4. Barriers to adopting platform security features This section explores why participants refrain from using PSFs and their perceptions of what such features offer (Section 5.4).
5. Opportunities for improvement and support This section explores participants’ wants and needs for improving current PSFs and vulnerability management practices in general (Section 5.5).
@@ 5 Results
In the following section, we report and discuss the results for 22 semi-structured interviews with open-source maintainers. In our reporting, we adhere to the structure of the interview guide described in Section 3.2. We report quotes as transcribed with minor grammatical corrections and omissions. e.g., using “[…]”, for readability.
@@
F# Factors and tooling Description RW L I Current practices F1 Email or mailing list Uses a private email or mailing list for vulnerabilities. [101, 68] 50 11 F2 Proactive disclosure Established disclosure process after a vulnerability patch. [101] 46 7 F3 Security policy Instructs users how to report vulnerabilities. [101] 37 10 F4 GitHub private reporting A built-in feature for privately reporting vulnerabilities. 20 9 F5 Automation tooling E.g., Dependabot, code scanning, secret scanning. [86, 35, 56, 81, 7, 79] 16 12 F6 GitHub Issues Vulnerabilities are reported publicly with GitHub Issues. [29] 9 6 F7 External tooling Utilizes tooling or reports outside of GitHub. [56, 101, 40] 3 2 F8 Ignores vulnerabilties Ignores vulnerability reports or does nothing in response. 2 General challenges F9 Supply chain trust E.g., waiting for upstream dependencies to deploy a fix. [9, 101, 31, 85, 45] 36 6 F10 Lack of understanding E.g., knowledge gaps, how to start developing a patch. [9, 13] 35 12 F11 Lack of time Balancing vulnerability priority with other OSS tasks. [67, 101, 13] 25 9 F12 Lack of resources E.g., “I do almost everything myself” (P4). [101, 25] 23 11 F13 CVE relationships Negative experiences with CVE-assigned vulnerabilities. 8 9 F14 Lack of procedure Maintainers are not sure how to deal with vulnerabilities. [101] 7 5 F15 Disclosure coordination How to reach out to all users and what to disclose. [9] 7 4 F16 Negative attitudes Past negative interactions with reporters or companies. [77, 88, 36, 63, 67, 34] 6 9 PSF challenges F17 Not enough automation More automation needed in PSF functionality. [66, 23] 39 11 F18 Too much noise E.g., too many false positives. [101, 100, 56, 79, 103, 102, 22, 105] 24 10 F19 Vulnerability scoring E.g., having a hard time calculating CVSS scores. 21 9 F20 CI processes are missing GitHub Actions & tests are not available in private forks. 8 7 F21 Broken tests & builds Failing builds & tests after vulnerability patch is applied. [105, 89, 79, 22] 6 7 F22 Cluttered notifications E.g., vulnerable dependencies presentation is cluttered. 4 4 F23 Too much manual setup E.g., manually enabling PSFs on all projects. 3 3 PSF barriers F24 Lack of awareness E.g., not knowing about private vulnerability reporting. [12] 33 12 F25 Complex to set up or use PSFs are too complex for maintainers with limited time. 26 7 F26 They are unnecessary Perception that using PSFs is not needed. 23 6 F27 Reputation concerns Past vulnerabilities reflect a negative project reputation. 11 4 F28 Bad UI presentation Hard to find or hidden, “second-class features” (P10). [32, 64, 40] 8 6 F29 Lack of motivation Feeling burned out from other OSS tasks. [88, 44, 74, 10, 67] 5 4 F30 Not sure what they do Unsure of PSFs’ functionality, benefits, etc. 4 2 Maintainer wants F31 Assisted triaging Automation for impact analysis, patch development, etc. [66, 23, 102] 39 9 F32 Assisted PSF setup More guidance for the PSF setup process. 37 10 F33 Security-specific funding Opportunities for funding OSS security efforts. [101] 32 7 F34 User-friendly resources Tailored for those without a security background. 17 8 F35 Gamification for projects E.g., a green shield for projects with proper PSFs set up. [82, 97, 31] 13 6 F36 PSF checklist To-do list for enabling easily configurable PSFs. 15 3 F37 Nudges or reminders Nudges reminding maintainers to consider using PSFs. 12 5
@@
Table 2: 37 factors and tooling codes, shortened to preserve space, identified from the listing survey study and interview study divided into five categories as a result of software vulnerability management efforts. Acronyms listed represent, respectively: platform security features (PSF), existing related work (RW), the number of listing (L) survey study mentions ($n_{1}=80$), and the number of interview (I) study mentions ($n_{2}=22$). The extended table without omitted codes can be found in our artifact [2].
@@ 5.1 Current practices used for vulnerability management
Emails and external tooling or reports (F1 & F7) Using a form of a security contact email or mailing list consisting of project maintainers was the most listed form of current vulnerability management practices ($L=50$, $I=11$). This is consistent with a finding from Wermke et al., where authors identified that having a security contact point was the commonly mentioned aspect of a security policy [101].
@@
Proactive disclosure process (F2) Having an established disclosure process after patching vulnerabilities consisting of project maintainers was the second-highest form of current vulnerability management practices ($L=46$, $I=7$). This demonstrates a sense of responsibility within the OSS community since it ensures affected dependent client projects within reach are notified to upgrade [81].
@@
Security policies (F3) The next current practice most listed is having a project security policy ($L=37,I=10$). OSS security policies have mostly been used to inform reporters how to properly communicate vulnerabilities to maintainers via security contact [101], but they are generally underused [11].
@@
Private vulnerability reporting (F4) Another GitHub platform security feature listing participants mentioned frequently is private vulnerability reporting ($L=20,I=9$). This particular feature allows contributors to privately report vulnerabilities within the GitHub platform. OSS maintainers can then review a report with functionalities to update its severity, invite others to develop a fix, and decide whether or not to request a CVE.
@@
Automation tooling (F5) Another method of current practices listing and interviewees mention is the usage of automated dependency analysis tooling and code scanning $(L=16,I=12)$. Dependency analysis provides alerts of library upgrades, e.g., patched versions, while code scanning can reveal vulnerabilities, e.g., secrets, as a result of running static analysis.
@@
“The first one that I use quite often is Renovate, that is a tool in Github easily available where you can configure: I want this and this upgraded like that, and you can have all kinds of settings and then it automatically gives you a pull request […] with a dependency update and automatically, the test pipeline fires.” (P1).
@@
GitHub Issues (F6) Some listing participants and interviewees indicate using GitHub issues for reporting vulnerabilities ($L=9,I=6$). When a GitHub Issue is submitted, it is publicly accessible in the Issues tab located at the top of the project landing page.
@@
Ignoring vulnerabilties (F8) Two listing participants mention how they ignore vulnerabilities. One attributed this behavior to a lack of motivation for setting up any security features or tooling and indicated that “financial support or sponsorship” would be most beneficial for them to improve their vulnerability management practices. The other respondent listed “unless the project is extremely sensitive - avoid implementing any policy” and expressed the perception that having security features enabled for projects is unnecessary. Though this respondent listed that they were interested in participating in our interview study, they did not respond with any contact information or reach out to us directly.
[…]
@@ 5.2 Challenges with vulnerability management
All interviewees described wearing many hats when it comes to OSS maintenance, some of who are lone maintainers ($I=11$). General vulnerability management challenges OSS maintainers face range from trusting the software supply chain to lack of time and resources to issues with CVEs. Participants also describe how a lack of standardized vulnerability handling procedures and implementing proper coordinated disclosure are prominent challenges. Some bring up the idea of having “imposter syndrome” when receiving vulnerability reports, while others share negative reporter experiences.
Supply chain trustworthiness (F9) Supply chain trust was the most listed challenge with vulnerability management ($L=36,I=6$). Use cases and duties related to the software supply chain in OSS include adopting upstream dependencies, tracking the status of dependencies, and updating dependencies in a timely manner, e.g., when a vulnerability is patched.
@@
Lack of understanding (F10) The second most listed challenge for conducting vulnerability management is a lack of understanding ($L=35,I=12$). Not being able to effectively understand reported vulnerabilities can lead to delays in patching and disclosure to affected dependent client projects.
@@
Lack of time and resources (F11 & F12) The next most listed software vulnerability management challenges are the lack of time ($L=25,I=9$) and resources ($L=23,I=11$). The time OSS maintainers spend on developing and managing projects is limited, e.g., hobbyists. Further, resources are scarce, e.g., large-scale software ecosystems like Python Package Index and NPM are primarily made up of projects with a single maintainer [20, 21].
@@
Negative relationships with CVEs (F13) Participants also expressed bad relationships with CVEs and the overall CVE process as a challenge when conducting vulnerability management ($L=8,I=9$). CVEs play an important role in the overall disclosure process and help promote transparency in the OSS ecosystem, e.g., upgrade notifications.
@@
Lack of procedures and coordinated disclosure (F14 & F15) Though established procedures can help alleviate the receive-to-resolve timeline, the lack of procedures is a challenge OSS maintainers also face ($L=7,I=5$). Further, participants also express challenges with being able to effectively reach affected dependent projects and users ($L=7,I=4$).
@@
Negative attitudes (F16) The least listed challenge of vulnerability management is negative attitudes from reporters ($L=6,I=9$). The OSS ecosystem is human-centered, e.g., via community and collaboration, and is not new to toxic interactions [88, 77]. One interviewee described experiences with companies demanding pentests. Participants mention the argumentative nature with vulnerability reporters ($L=3,I=7$) and those with sole intentions to make money ($L=3,I=5$).